OpenStudy (anonymous):

expand the following function ln(a(b-4)^2 e^(4c^2))

OpenStudy (freckles):

\[\ln(a(b-4)^2e^{4c^2})\] it this right:

OpenStudy (anonymous):

yes

OpenStudy (freckles):

have you tried using the product rule yet? \[\ln(a \cdot b)=\ln(a)+\ln(b)\]

OpenStudy (anonymous):

this is what i got but im not sure lna+2ln(b-4)+4c^2

OpenStudy (anonymous):

this is what i got but im not sure lna+2ln(b-4)+4c^2

OpenStudy (freckles):

\[\ln(a)+\ln((b-4)^2)+\ln(e^{4c^2} ) \text{ by product rule } \\ \ln(a)+2 \ln(b-4)+\ln(e^{4c^2}) \text{ by power rule } \\ \ln(a)+2 \ln(b-4)+4c^2 \text{ using the inverse fact}\] you are right!

OpenStudy (anonymous):

thanks